Evaluate
(z+y)/6 for z = 33 and y = 15
(z+y)/6=

Answer by fcabanski(1391)

Substitute the given values for z and y into the equation.


(z+y)/6 when z=33 and y=15 is (33+15)/6 = 48/6 = 8

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com